# IllusionVQA: A Challenging Optical Illusion Dataset for Vision Language Models

### v2

@@ 5.2 IllusionVQA-Soft-Localization
Table 3: Performance of VLMs on IllusionVQA-Soft-Localization. Human performance denotes the majority-vote performance of three human evaluators on 200 randomly picked instances.
Table 3 shows that the localization of optical illusions is a significant challenge for VLMs but quite straightforward for humans. Small open-source VLMs, particularly, show close to random performance (25%). GPTV and Gemini-Pro fare somewhat better but fall short of human performance. On 200 randomly selected examples, all three human evaluators achieved perfect accuracy. Even with ICL and CoT, the performance of GPT4V in the IllusionVQA-Soft-Localization dataset remains subpar, achieving only 49.7% accuracy  [^3] Random guessing would yield a 25% accuracy rate.. This unexpectedly poor performance motivated us to conduct additional experiments in Appendix F.
@@
We created a new dataset for soft localization using different brands of sports cars and asked the VLM “Which car is a [BRAND]?" with the same four options as in Figure 5). This task remains moderately challenging for VLM since sports cars of different brands share the same basic shape and differ in details. Although smaller VLMs struggled with this task, both GPT4V and Gemini-Pro achieved 100% accuracy. GPT4V and Gemini’s stellar performance in soft localization of ordinary objects makes their unexpectedly low performance on IllusionVQA-Soft Localization surprising. This is likely because differentiating geometric illusions from ordinary objects requires spatial reasoning (a known weakness of VLMs OpenAI (2023)) instead of feature extraction. We present the details regarding the dataset and the model performance in Appendix F.
@@
In-context learning (ICL) Brown et al. (2020) and Chain-of-Thought (CoT) Wei et al. (2022) are widely used techniques that boost LLM performance. Research has shown that these techniques are effective in vision language tasks Yang et al. (2023), but are still understudied. We find that these methods are highly model-dependent.
Table 3 shows that Gemini-Pro performs better than GPT4V (+3.5%) when evaluated without ICL or CoT. However, GPT4V shows the highest accuracy overall when evaluated with 4-shot+CoT (+9.7% gain) while Gemini-Pro’s performance degrades substantially (-9.6%). We hypothesize this is because GPT4V is likely the larger model and exhibits stronger emergent ICL abilities Wei et al. (2023). As illustrated in Figure 6, there are a substantial number of instances in which the 0-shot evaluation yielded the correct answer, while 4-shot+CoT reasoning failed to do so. This implies that ICL and CoT are not universally applicable to IllusionVQA-Soft-Localization even with GPT4V.
(a) GPT4 - correctly ans.
(b) GPT4 - incorrectly ans.
(c) Gemini-Pro - correctly ans.
(d) Gemini-Pro - incorrectly ans.
@@
To probe the weakness of VLMs in soft localization, we conducted additional experiments in which we asked GPT4V and Gemini-Pro with 4 examples (4 shots) and presented the same examples as questions. In this case, the exact image, the question, and the correct answer are in the context window of the VLM. Surprisingly, neither GPT4V nor Gemini-Pro can successfully answer all questions. This highlights a puzzling limitation of ICL for VLMs where models fail even when the correct answer is in the context which points to over-dependence on language priors Goyal et al. (2017); Jabri et al. (2016) and visual information is all but ignored. This phenomenon persists in all the combinations we tried. We provide the prompt and result for one such combination in Appendix H.
@@ 6.1 Illusion-Resistant VLMs for Robotics
Vision Language Models (VLMs) are increasingly being deployed in real-world applications, particularly in robotics Zitkovich et al. (2023); Wake et al. (2023), where they play a crucial role in enabling robots to navigate and interact with their environment. Embodied AI systems need to demonstrate consistent performance across a wide range of real-world situations, including those involving deceptive designs, optical illusions, obscured objects, or non-traditional image orientations.
@@ 6.2 Thinking Fast and Slow - Comparing VLM and Human Response Times
Kahneman (2011) popularized the concept of two modes of thinking: ‘System 1’, which is fast and instinctive, and ‘System 2’, which is slower, more deliberate, and logical. Due to their autoregressive architecture, current state-of-the-art LLMs are primarily capable of ‘System 1’ thinking Bubeck et al. (2023). Approximating ‘System 2’ reasoning with LLMs has motivated significant research efforts such as Chain-of-Thought Wei et al. (2022), Tree-of-Thought Yao et al. (2024), and other related methods.
@@ 7 Conclusion
We demonstrate that current state-of-the-art VLMs struggle with understanding optical illusions and largely falter in identifying geometrically impossible objects within a scene. We find that GPT4V maintains substantial leads in most types of illusion and that there is a noticeable performance gap between open-source and closed-source VLMs. We validate the effectiveness of few-shot learning for GPT4V and Gemini-Pro in the comprehension task but show that few-shot learning and Chain-of-Thought reasoning are not universally effective in the localization task. Additionally, we discover that GPT4V and Gemini-Pro fail to correctly answer illusion localization questions even when the correct answer is provided in the context. We release our dataset and evaluation code and hope to stimulate further research in bridging the gap between human and machine performance in challenging vision-related tasks.
Limitation
@@
    Evaluation Scope: Due to the OpenAI API rate limits (500 requests/day), we were unable to evaluate GPT4V in multiple evaluation runs. It is possible that more advanced evaluation techniques Lei et al. (2024) or task-specific fine-tuning of open-source VLMs with sufficient data might yield improvements on IllusionVQA-Soft-Localization.
@@ Acknowledgement
We express our gratitude to the contributors and the owner of moillusions.com for curating and hosting a collection of high-quality optical illusions online. We thank Michael Bach for allowing us to gather optical illusions and their descriptions from his excellent repository. Lastly, we appreciate the efforts of our human evaluators who diligently evaluated the IllusionVQA dataset.
@@ Appendix A Preliminaries
An optical illusion refers to a visual perception phenomenon in which the visual system misinterprets reality, resulting in a disparity between what is perceived and what exists. Researchers have proposed that optical illusions mainly arise from discrepancies in a person’s perceptual knowledge and conceptual knowledge Gregory (1997b); Carbon (2014). As a concrete example, our fast perceptual knowledge perceives Figure 3(a) as an elephant, yet upon closer inspection, we understand that the diagram does not align with our preexisting concepts regarding geometry.
Figure 7: Left: Kitaoka’s ‘Throwing cast nets’ Conway et al. (2005), an example of cognitive impenetrability. Even though we know it is a static image, the illusion of motion persists. Right: Schroeder stairs Rosenholtz (2011), an example of an ambiguous or bistable image. The drawing may be perceived as either a staircase leading from left to right downward or as the same staircase only turned upside down.
@@
is a property of the basic information processes of cognition. If a process is cognitively impenetrable, then the operation of this process cannot be altered by changing the contents of mental representations (i.e., an agent’s beliefs, desires, or goals). This is because the process is “wired in" and its operation must be explained by appealing to neuroscience Dawson (2017). Select visual illusions (Figure 7 - Left) exhibit this phenomenon because the illusion persists even if we know or are made aware of the illusion.
@@
are images that create ambiguity between two or more distinct image forms. Classic examples of ambiguous images include Rubin’s Vase Hasson et al. (2001), Schroeder stairs (Figure 7 - Right), and Necker’s Cube Rosenholtz (2011) among many others.
@@ Appendix B Object Localization with VLMs
Object DetectionGirshick et al. (2014) is a core computer vision task in which models output the bounding box coordinates of objects in an image. Although pure vision models such as DETRZong et al. (2023) are state of the art in object detection, there has been considerable progress in object detection using pretrained VLMs Wang et al. (2023); Bai et al. (2023). However, most VLMs do not natively support object detection and perform significantly worse than vision-only models. Consequently, we opt for a simpler task, soft localization, where the VLMs are prompted to output the approximate position of objects (i.e. left side or right side) instead of exact bounding box coordinates.
@@ Appendix C Literature Review of Illusion Classification
Classifying optical illusions has posed significant challenges, with no definitive consensus reached within the scientific community. We review two prominent categorization approaches proposed by Gregory (1997b) and Bach and Poloschek (2006) as precursors to our categorization of IllusionVQA.
Gregory (1997b) proposed a four-fold categorization of optical illusions based on their underlying causes: physical, physiological, knowledge-based, and rule-based. Physical illusions, like those induced by fog or lens distortions, directly interfere with image formation. Physiological illusions, such as scintillations from pressure on the eye, arise from disruptions within the visual system. Knowledge-based illusions, exemplified by interpreting 2D sketches as 3D objects, leverage prior knowledge to misinterpret sensory input. Finally, rule-based illusions, such as Kanizsa triangles Kanizsa (1976), exploit the visual system’s tendency to complete patterns or infer missing information, leading to illusory percepts. Gregory (1997b) proposes a complementary classification of optical illusions based on appearance and rooted in natural language, namely ambiguities (Figure 7), distortions (Figure 3(b)), paradoxes (Figure 3(a)), and fictions (Figure 3(c)). We adopt this classification scheme for IllusionVQA in section 3.3.
@@ Appendix D IllusionVQA-Comprehension Categories and Question Types
Impossible Object 134 Objects that cannot exist in 3D space. Penrose Triangle Penrose and Penrose (1958)
@@
Size 46 Two objects with the same dimensions look different or vice versa. Müller-Lyer illusion Judd (1905)
@@
Deceptive Design 37 Items have been painted to appear as something else. Trompe-l’œil art Wade and Hughes (1999)
Angle Illusion 26 Parallel lines look curved or angled, and vice versa. Zöllner illusion Zöllner (1860).
Color 23 Different colors or shades appear the same, and vice versa. Checker Shadow Illusion Adelson (1995)
@@
Positive-Negative Space 7 Images that create ambiguity between two or more distinct forms represented by two or more colors. Rubin’s vase Hasson et al. (2001)
Circle-Spiral 6 Circles appear as spirals due to peculiar color or background pattern. Fraser Spiral Fraser (1908)
@@ Appendix G Human Evaluation
All three evaluators were presented with the questions in the same order. Table 7 shows that the individual accuracy of the evaluators is comparable. The pairwise Cohen’s Kappa between the evaluators are 0.808 0.796 and 0.773, indicating a substantial level of inter-evaluator agreement  [^5] To determine the agreement metric, we considered each question-option pair as a singular question with a yes/no answer which denotes whether an evaluator has selected that option as the answer or not.. On IllusionVQA-Soft-Localization, all evaluators achieved perfect accuracy on the selected subset of 200 instances and the average response time was 5.68 seconds.

### v3

@@ 5.2 IllusionVQA-Soft-Localization
Table 3 shows that the localization of optical illusions is a significant challenge for VLMs but quite straightforward for humans. Small open-source VLMs, particularly, show close to random performance (25%). GPTV and Gemini-Pro fare somewhat better but fall short of human performance. On 200 randomly selected examples, all three human evaluators achieved perfect accuracy. Even with ICL and CoT, the performance of GPT4V in the IllusionVQA-Soft-Localization dataset remains subpar, achieving only 49.7% accuracy [^3] Random guessing would yield a 25% accuracy rate.. This unexpectedly poor performance motivated us to conduct additional experiments, presented in Appendix F.
@@
Table 3: Performance of VLMs on IllusionVQA-Soft-Localization. Human performance denotes the majority-vote performance of three human evaluators on 200 random instances.
@@
We created a new dataset for soft localization using different brands of sports cars and asked the VLM “Which car is a [BRAND]?" with the same four options as in Figure 4). This task remains moderately challenging for VLM since sports cars of different brands share the same basic shape and differ in details. Although smaller VLMs struggled with this task, both GPT4V and Gemini-Pro achieved 100% accuracy. GPT4V and Gemini’s stellar performance in soft localization of ordinary objects makes their unexpectedly low performance on IllusionVQA-Soft Localization surprising. This is likely because differentiating geometric illusions from ordinary objects requires spatial reasoning (a known weakness of VLMs (OpenAI, 2023)) instead of feature extraction. We present the details regarding the dataset and the model performance in Appendix F.
@@
In-context learning (ICL) (Brown et al., 2020) and Chain-of-Thought (CoT) (Wei et al., 2022) are widely used techniques that boost LLM performance. Research has shown that these techniques are effective in vision language tasks (Yang et al., 2023), but are still understudied.
Table 3 shows that Gemini-Pro performs better than GPT4V (+3.5%) when evaluated without ICL or CoT. However, GPT4V shows the highest accuracy overall when evaluated with 4-shot+CoT (+9.7% gain) while Gemini-Pro’s performance degrades substantially (-9.6%). We hypothesize this is because GPT4V is likely the larger model and exhibits stronger emergent ICL abilities (Wei et al., 2023). As illustrated in Figure 5, there are a substantial number of instances in which the 0-shot evaluation yielded the correct answer, while 4-shot+CoT reasoning failed to do so. This implies that ICL and CoT are not universally applicable to IllusionVQA-Soft-Localization.
(a) GPT4
- correctly ans.
(b) GPT4
- incorrectly ans.
(c) Gemini-Pro
- correctly ans.
(d) Gemini-Pro
- incorrectly ans.
@@
To probe the weakness of VLMs, we conducted additional experiments in which we asked GPT4V and Gemini-Pro with 4 examples (4 shots) and presented the same examples as questions. In this case, the exact image, the question, and the correct answer are in the context window of the VLM. Surprisingly, neither GPT4V nor Gemini-Pro can successfully answer all questions. This highlights a puzzling limitation of ICL for VLMs where models fail even when the correct answer is in the context which points to over-dependence on language priors (Goyal et al., 2017; Jabri et al., 2016) and visual information is all but ignored. This phenomenon persists in all the combinations we tried. We provide the prompt and result for one such combination in Appendix H.
@@ 6.1 Illusion-Resistant VLMs for Robotics
Vision Language Models (VLMs) are increasingly being deployed in real-world applications, particularly in robotics (Zitkovich et al., 2023; Wake et al., 2023), where they play a crucial role in enabling robots to navigate and interact with their environment. Embodied AI needs to demonstrate consistent performance across a wide range of real-world situations, including those involving deceptive designs, optical illusions, and obscured objects.
@@ 6.2 Thinking Fast and Slow - Comparing VLM and Human Response Times
Kahneman (2011) popularized the concept of two modes of thinking: ‘System 1’, which is fast and instinctive, and ‘System 2’, which is slower, more deliberate, and logical. Due to their autoregressive architecture, current state-of-the-art LLMs are primarily capable of ‘System 1’ thinking (Bubeck et al., 2023). Approximating ‘System 2’ reasoning with LLMs has motivated significant research efforts such as Chain-of-Thought (Wei et al., 2022), Tree-of-Thought (Yao et al., 2024), and other related methods.
@@ 7 Conclusion
We demonstrate that current state-of-the-art VLMs struggle with understanding optical illusions and largely falter in identifying geometrically impossible objects. We find that GPT4V maintains substantial leads in most illusion types and that there is a noticeable performance gap between open-source and closed-source VLMs. We validate the effectiveness of few-shot learning for GPT4V and Gemini-Pro in the comprehension task but show that few-shot learning and Chain-of-Thought reasoning are not universally effective in the localization task. Additionally, we discover that GPT4V and Gemini-Pro fail to correctly answer illusion localization questions even when the correct answer is provided in the context. We release our dataset and evaluation code to stimulate further research.
Limitations
@@
    Evaluation Scope: Due to the OpenAI API rate limits (500 requests/day), we were unable to evaluate GPT4V in multiple evaluation runs. It is possible that more advanced evaluation techniques (Lei et al., 2024) or task-specific fine-tuning of open-source VLMs with sufficient data might yield improvements on IllusionVQA-Soft-Localization.
@@ Acknowledgement
We thank the contributors and the owner of moillusions.com for curating and hosting a collection of high-quality optical illusions online. We thank Michael Bach for allowing us to gather optical illusions and their descriptions from his repository. Lastly, we appreciate the efforts of our human evaluators who diligently evaluated the IllusionVQA dataset.
@@ Appendix A Preliminaries
An optical illusion refers to a visual perception phenomenon in which the visual system misinterprets reality, resulting in a disparity between what is perceived and what exists. Researchers have proposed that optical illusions mainly arise from discrepancies in a person’s perceptual knowledge and conceptual knowledge (Gregory, 1997b; Carbon, 2014). As a concrete example, our fast perceptual knowledge perceives Figure 2(a) as an elephant, yet upon closer inspection, we understand that the diagram does not align with our preexisting concepts regarding geometry.
Figure 6: Left: Kitaoka’s ‘Throwing cast nets’ (Conway et al., 2005), an example of cognitive impenetrability. Even though we know it is a static image, the illusion of motion persists. Right: Schroeder stairs (Rosenholtz, 2011), an example of an ambiguous or bistable image. The drawing may be perceived as either a staircase leading from left to right downward or as the same staircase only turned upside down.
@@
refers to a property of fundamental cognitive information processes. When a process is cognitively impenetrable, its operation remains unaffected by changes in the contents of mental representations, such as an agent’s beliefs, desires, or goals. This is because the process is “wired in" and can only be explained through neuroscience (Dawson, 2017). The visual illusions shown in Figure 6 (Left) illustrate this phenomenon, as the illusion persists even when we are aware of it.
@@
are images that generate ambiguity between two or more distinct forms. Classic examples of such images include Rubin’s Vase (Hasson et al., 2001), Schroeder stairs (Figure 6 - Right), and Necker’s Cube (Rosenholtz, 2011), among others.
@@ Appendix B Object Localization with VLMs
Object Detection (Girshick et al., 2014) is a core computer vision task in which models output the bounding box coordinates of objects in an image. Although pure vision models such as DETR(Zong et al., 2023) are state of the art in object detection, there has been considerable progress in object detection using pretrained VLMs (Wang et al., 2023; Bai et al., 2023). However, most VLMs do not natively support object detection and perform significantly worse than vision-only models. Consequently, we opt for a simpler task, soft localization, where the VLMs are prompted to output the approximate position of objects (i.e. left side or right side) instead of exact bounding box coordinates.
@@ Appendix C Literature Review of Illusion Classification
Classifying optical illusions has posed significant challenges, with no definitive consensus reached within the scientific community. We review two prominent categorization approaches proposed by Gregory (1997b) and Bach & Poloschek (2006) as precursors to our categorization of IllusionVQA.
Gregory (1997b) proposed a four-fold categorization of optical illusions based on their underlying causes: physical, physiological, knowledge-based, and rule-based. Physical illusions, like those induced by fog or lens distortions, directly interfere with image formation. Physiological illusions, such as scintillations from pressure on the eye, arise from disruptions within the visual system. Knowledge-based illusions, exemplified by interpreting 2D sketches as 3D objects, leverage prior knowledge to misinterpret sensory input. Finally, rule-based illusions, such as Kanizsa triangles (Kanizsa, 1976), exploit the visual system’s tendency to complete patterns or infer missing information, leading to illusory percepts. Gregory (1997b) proposes a complementary classification of optical illusions based on appearance and rooted in natural language, namely ambiguities (Figure 6), distortions (Figure 2(b)), paradoxes (Figure 2(a)), and fictions (Figure 2(c)). We adopt this classification scheme for IllusionVQA in section 3.3.
@@ Appendix D IllusionVQA-Comprehension Categories and Question Types
Impossible Object 134 Objects that cannot exist in 3D space. Penrose Triangle (Penrose & Penrose, 1958)
@@
Size 46 Two objects with the same dimensions look different or vice versa. Müller-Lyer illusion (Judd, 1905)
@@
Deceptive Design 37 Items have been painted to appear as something else. Trompe-l’œil art (Wade & Hughes, 1999)
Angle Illusion 26 Parallel lines look curved or angled, and vice versa. Zöllner illusion (Zöllner, 1860).
Color 23 Different colors or shades appear the same, and vice versa. Checker Shadow Illusion (Adelson, 1995)
@@
Positive-Negative Space 7 Images that create ambiguity between two or more distinct forms represented by two or more colors. Rubin’s vase (Hasson et al., 2001)
Circle-Spiral 6 Circles appear as spirals due to peculiar color or background pattern. Fraser Spiral (Fraser, 1908)
@@ Appendix G Human Evaluation
All three evaluators were presented with the questions in the same order. As shown in Table 7, their individual accuracy levels are comparable. The pairwise Cohen’s Kappa values between the evaluators are 0.808, 0.796, and 0.773, indicating a substantial level of inter-evaluator agreement. [^5] For the agreement metric, we treated each question-option pair as a distinct question with a yes/no answer, indicating whether an evaluator selected that option as the correct answer. On the IllusionVQA-Soft-Localization, all evaluators achieved perfect accuracy across 200 selected instances, with an average response time of 5.68 seconds.
@@ Appendix I Examples of Images Removed During Preprocessing
Appendix J Updated IllusionVQA-Comprehension Results
Class # 0-shot 4-shot Human
InternVL2
8B
Pali
Gemma
Gemini
1.0-Pro
Claude 3.5
Sonnet
GPT4V GPT4o
Gemini
1.0-Pro
Claude 3.5
Sonnet
GPT4V GPT4o
Impossible Object 134 49.25 32.09 56.72 64.93 55.22 63.43 56.72 63.43 58.96 61.94 98.51
Real-Scene 64 40.63 35.94 46.88 54.69 57.81 64.06 46.88 57.81 54.69 57.81 98.44
Size 46 43.48 15.22 45.65 50.00 58.70 45.65 52.17 80.43 69.57 93.47 63.04
Hidden 45 44.44 33.33 42.22 37.78 51.11 66.67 48.89 44.44 46.67 48.89 100
Deceptive Design 37 37.84 32.43 64.86 70.27 70.27 72.97 67.57 67.57 72.97 78.38 94.59
Angle Illusion 26 50.00 26.92 53.85 73.08 69.23 50.00 50.00 73.08 84.62 80.77 84.62
Color 23 26.09 34.78 17.39 65.22 69.57 52.17 17.39 86.96 82.61 78.26 60.87
Edited-Scene 21 66.67 42.86 66.67 61.90 71.43 80.95 66.67 71.43 80.95 85.71 100
Upside-Down 7 85.71 42.86 57.14 85.71 71.43 71.43 57.14 85.71 71.43 42.86 100
Pos.-Neg. Space 7 42.86 42.86 85.71 71.43 57.14 85.71 71.43 71.43 85.71 71.43 100
Circle-Spiral 6 0.00 0.00 33.33 33.33 50.00 50.00 33.33 50.00 33.33 50.00 66.67
Miscellaneous 19 42.11 31.58 52.63 47.37 42.11 52.63 57.89 52.63 42.11 52.63 89.47
Total 435 45.06 31.26 51.26 59.08 58.85 62.53 52.87 66.44 62.99 67.12 91.03
Table 8: IllusionVQA-Comprehension results updated to include state-of-the-art VLMs as of July 2024. Added evaluation of InternVLM2 (Chen et al., 2024), PaliGemma-3B (Beyer et al., 2024), GPT-4o (OpenAI, 2024) and Claude 3.5 Sonnet Anthropic (2024). Categories where accuracy has improved using 4-shot prompting are underlined.